OpenStudy (anonymous):

if hx=x2−5, find h (-3) A. -14 B. 15 C. -11 D. 4

OpenStudy (anonymous):

Okay, we have -3=x2-5. Do you know what to do next?

OpenStudy (anonymous):

i got x=-14

OpenStudy (anonymous):

Hmm. I got a different answer. Lets check my math. So we had -3=x2-5. We need to isolate the X correct?

OpenStudy (anonymous):

x=1.414214,−1.414214

OpenStudy (anonymous):

to get rid of the � replace it all with : -